# ComplexityMeasures.jl: scalable software to unify and accelerate entropy and complexity timeseries analysis

### v1

@@ Abstract
In the nonlinear timeseries analysis literature, countless quantities have been presented as new “entropy” or “complexity” measures, often with similar roles. The ever-increasing pool of such measures makes creating a sustainable and all-encompassing software for them difficult both conceptually and pragmatically. Such a software however would be an important tool that can aid researchers make an informed decision of which measure to use and for which application, as well as accelerate novel research. Here we present ComplexityMeasures.jl, an easily extendable and highly performant open-source software that implements a vast selection of complexity measures. The software provides 1530 measures with 3,834 lines of source code, averaging only 2.5 lines of code per exported quantity (version 3.5). This is made possible by its mathematically rigorous composable design. In this paper we discuss the software design and demonstrate how it can accelerate complexity-related research in the future. We carefully compare it with alternative software and conclude that ComplexityMeasures.jl outclasses the alternatives in several objective aspects of comparison, such as computational performance, overall amount of measures, reliability, and extendability. ComplexityMeasures.jl is also a component of the DynamicalSystems.jl library for nonlinear dynamics and nonlinear timeseries analysis and follows open source development practices for creating a sustainable community of developers.
@@ 1 Introduction
A large aspect of nonlinear timeseries analysis [1, 2, 3] is concerned with extracting quantities (i.e. computing various statistics) from timeseries that quantify some property of the underlying dynamics that generated the timeseries. The purpose of these statistics can be to distinguish one type of dynamics from another [4], to classify timeseries into classes with different dynamics [5, 6], to quantify directional associations between time series  [7], which in turn can be integrated into frameworks for conditional independence testing between time series  [8], and more. Most of these statistics are labelled complexity measures, because they quantify in some way the amount of complexity in the system. Although the word “complex” does not have a widely-accepted definition yet [9], it typically describes something that is both not regular nor purely stochastic.
The majority of complexity measures are based on some form of axiomatically well-founded entropy. For example, the permutation entropy [10] and the wavelet entropy [11] are based on the Shannon entropy (Eq. 1). Other complexity measures are not entropies in the formal mathematical sense, but are inspired by, or related to, entropies. Approximate entropy [12] and sample entropy [13], for example, are entropy rates, rather than entropies. Like an entropy, these entropy-like measures will typically yield higher numerical values for more “complex” data, where ”complex” has a measure-specific definition. In the rest of the text we will be referring to all these entropy or complexity quantities simply as complexity measures, and only use the word “entropy” if its rigorous mathematical definition is important in the context.
@@ 1.1 Example of estimating a complexity measure
There exists a surprisingly large amount of complexity measures in the literature. For simplicity, we will start by focusing on the largest class of complexity measures: those that are functionals of probabilities mass functions (PMFs). Here, we will deal exclusively with empirical PMFs, which are PMFs estimated from data. We will use the terms "probabilities" and empirical PMFs interchangeably.
All discrete probability-based complexity measures apply the same fundamental steps, and can unified under one estimation pipeline. To estimate probabilities from observed data, it is necessary to first define a specific and countable outcome space $\Omega$. The goal is then to assign a probability $p_{i}=p(\omega_{i})$ to each outcome $\omega_{i}\in\Omega$ based on the input data. To do so, the input data needs to be mapped (encoded, or discretized) onto the elements of $\Omega$. Sometimes, this procedure is also called symbolization. After encoding, we can form an empirical distribution over the encoded symbols. An empirical distribution in this context just means a histogram, or a normalized pseudo-histogram, depending on the structure of the outcome space. Next, the probabilities $p(\omega_{i})$ are estimated based from this empirical distribution, for example using relative frequency estimation. Finally, once a probability vector $\mathbf{p}=\{p(\omega_{i})\}_{i=1}^{N}$ has been constructed, these probabilities can be given as an input to some probabilities functional.
As an example, let’s say we’d want to compute the order-3 permutation entropy [10] for an input time series $x$. To do so, we first construct a 3-dimensional embedding of $x$. Three-element state vectors can be ordered in $3!=6$ different ways. We’ll consider each one of these possible orderings, which are also called ordinal patterns, as separate outcomes. We can then proceed by mapping each state vector in the embedding uniquely onto one of the ordinal patterns  [^1] Assuming there are no ties in the state vector., which is the same as saying that we encode the input data. To estimate probabilities, we can then simply count the relative frequency/occurrence of the different ordinal patterns and normalize these counts to sum to 1 (also called plug-in, or maximum likelihood estimation). Finally, these probabilities are given to the Shannon entropy formula [14]
@@
This estimator of the Shannon entropy is called the naive, or plug-in estimator, and returns a non-negative number which is indicative of the "complexity" of the input data.
@@ 1.2 Combinatorial explosion of complexity measures
In the first step of the procedure outlined above, we implicitly chose an outcome space (a way to map data into outcomes). We could decide to use any other outcome space instead. One commonly used class of outcome spaces are rectangular binnings (i.e., histograms), in which each each data point is mapped onto one bin according to its value. Other examples of outcome spaces are dispersion patterns [15], binned cosine similarities [16], binned state vector distances [17], sorting complexity [18], which all consist of an initial embedding step, after which the embedding vectors are encoded using some procedure that cleverly highlights some interesting property of the underlying data.
@@
In the third step, we could have instead considered any of the other theoretically well-founded information measures, for example Rényi entropy ($H_{R}$) [22], Tsallis entropy ($H_{T}$) [23], Kaniadakis entropy ($H_{K}$) [24], Curado entropy ($H_{C}$) [25] and the Anteneodo-Plastino streched exponential entropy ($H_{AP}$) [26], the lesser known Shannon extropy ($J_{S}$) [27], Rényi extropy [28] or Tsallis extropy ($J_{T}$) [29], or any other probabilities functional that in some way quantify complexity.
@@
Thus, excluding parameterizations, there are four degrees of freedom when computing a discrete, probabilities-based complexity measure: the discretization/encoding procedure, the probabilities estimator, the information measure, and the estimator for the information measure. Now let’s assume that the literature describes $N_{O}$ different outcome spaces, $N_{P}$ different ways of estimating probabilities, and $N_{C}$ different probabilities-based complexity measures. Assume the number of estimators for a particular measure $M_{i}$ is $N_{M_{i}}$. Then the number of total computable PMF-based complexity measures is
@@
We quickly realize that there is a vast set possible complexity measures, varying across all four degrees of freedom, with differences ranging from very minor technicalities to major conceptual differences, yet all quantifying complexity in some unique way. Adding (for example) one more entropy definition $N_{C}\to N_{C}+1$ drastically increases the total amount of computable measures, because there are $N_{O}\times N_{P}$ potential ways of computing it from data for every unique estimator of this new measure. We call this the combinatorial explosion of complexity measures. In this context, the traditional software design approach of implementing one function for each measure is not scalable as it requires adding many more functions when wanting to add “only” one more complexity measure definition.
The large number of possible complexity measures could provide great opportunities for future research into complexity quantification. However, a systematic and easy-to-use software for exploring and comparing these different measures, that is also easy to extend with new ones, and has the computational capacity to compute all of them quickly, is lacking.
@@
ComplexityMeasures.jl was built to fill this gap. It resolves the explosion problem by taking a fundamentally different approach: the software allows for composable instructions for how to compute a complexity measure. For discrete, probabilities-based measures, this entails composing instructions on which definition to use, which estimator to use, with which probabilities estimator to estimate probabilities, and which outcome space to use for the discretization (Figure 1). A similar approach is taken when estimating other complexity measures, which are described later in the paper. In practice this leads to an incredibly lean source code base that is also easily extendable. This new design approach avoids the one-function-per-estimator typical software design that leads to an unnecessarily large, hard-to-maintain code base that is also hard to bugfix efficiently or future-proof. Additionally, by developing ComplexityMeasures.jl following highest standards in scientific software development [38], we made it both easily extendable and highly performant, so that it is sufficiently future-proof and can handle all complexity measures present and to-be-added in the future.
In Sec. 2 we expose the design behind ComplexityMeasures.jl and how it integrates into a wider ecosystem enabling further innovation (such as CausalityTools.jl  [39]); in Sec. 3 we provide some selected examples that highlight how accessible yet powerful the library is; lastly in Sec. 4 we compare the software with existing alternatives in a comprehensive tabular format and show that across many objective aspects of comparison ComplexityMeasures.jl performs best.
@@ 2.1 Core functions
The design of ComplexityMeasures.jl is displayed in Fig. 1 and perfectly parallelizes the mathematically rigorous formulation of an information measure as we described it in Sec. 1.1. Software usage revolves around two functions: information and complexity. The first is called as information(discr_ent_est, prob_est, ospace, data), and estimates an information measure given the information measure estimator from a PMF, a probabilities estimator to map discrete outcomes into a PMF, and an outcome space to map input data into discrete outcomes. To estimate the permutation entropy (as originally defined in [10]) one would call information(Shannon(), RelativeAmount(), OrdinalPatterns(m = 3), input). For convenience and conciseness, simpler “shortcuts” are also possible in ComplexityMeasures.jl. For example, entropy(OrdinalPatterns(m = 3), data) uses default values for the probability estimator (RelativeAmount()) and the information measure estimator (Shannon()). For a few handpicked measures even simpler syntax is available, such as the call entropy_permutation(data; m = 3), which is equivalent with the previous ones. More syntax shortcuts are discussed in the software documentation.
Fig 1: The two central functions of ComplexityMeasures.jl and the inputs they expect.
For complexity measures that are not explicit functionals of a PMF, we have implemented a simpler design shown in Fig. 1b. There a function complexity takes as an input a “complexity estimator” the defines a complexity measure as well as includes instructions on how to compute it directly from input data. Differential (or continuous) information measure estimators follow the same design in the current software version. In the future we plan to make a dedicated probability density estimation interface for continuous measures, similar to our discrete estimation interface based on probability mass functions. Both information and complexity have a normalized option that returns the complexity measure divided by its maximum possible value. Similarly, a multiscale function computes the multiscale variant of a chosen complexity measure (a framework originally introduced in Ref. [40] as “multiscale sample entropy”). Both normalized and multiscale forms are valid for any complexity measure in the library since these forms are also based on a composable design.
@@
    Orthogonality of inputs (information estimator, probabilities, and outcomes). This means that if, e.g., a combination of a probability estimator works for a given outcome space, it is guaranteed to work for any outcome space implemented in the software. This guarantee occurs automatically due to the design (and also permitted by the Julia language multiple dispatch system [41]) and does not need to be enforced by the user or even the developer.
@@
    The software is easy to maintain. Finding a bug in e.g., creating a histogram of some (optionally symbolized) data requires fixing this bug in only a single function that does the histogram counting, not in potentially hundreds of functions that utilize histograms in some way.
@@
    Simple and scalable extensions. Adding a new outcome space requires writing code for one new type (Julia’s version of “classes”) and one mandatory function extension instructing how to discretize data for this outcome space. This is can be as simple as 10 lines of code. Yet, once implemented, this outcome space would allow the user, without writing any additional code, to compute any probabilities-based complexity measure using this outcome space, in combination with any probabilities estimation, any information measure, and any estimator of this measure, including even some non-information complexity measures such as missing patterns [42].
@@
    The probabilities themselves are directly accessible by the user and compose a comprehensive interface that we expand more in Sec. 2.2. Analyzing the probabilities directly may expose something interesting about the data that is “integrated away” by the complexity measure computation. For example, if one computes the associated probabilities for the 6 order-3 ordinal patterns of a timeseries coming for a logistic map, 2 of the 6 patterns will have 0 probability due to the logistic map’s dynamics. Information like this can only be obtained by looking at the probability mass function directly, which is hidden away in the majority of alternative software. Additionally, having access to the probabilities directly allows expanding and creating new complexity measures not published before like the example we showcase in Sec. 3.2.
@@ 2.2 Outcome spaces and probabilities
During the design of ComplexityMeasures.jl, it became obvious that the majority of complexity measures are estimated based on some PMF extracted from data. To create the extendable design mentioned above, behind the main functions information, complexity there stands a fully-fledged API (application programming interface) for extracting probabilities from data, based on the mathematically rigorous formulation of an outcome space. To the best of our knowledge, there isn’t any open source software in any programming language that provides such an extensive interface for extracting probabilities from data.
In ComplexityMeasures.jl we define an abstract hierarchy of types called OutcomeSpace. An instance o of a concrete implementation of an OutcomeSpace describes how to discretize data into discrete outcomes. Given o, we define many functions for handling probabilities. For example, total_outcomes returns the cardinality of o, while missing_outcomes returns the number of outcomes defined by o as possible but not present in the data, and counts returns a vector of integers, counting how many times was each outcome present in the input data. We further separate this API into outcome spaces that are counting-based, which allow mapping each element of input data into an integer, and non-counting based which cannot do this. More details on this can be found on the developer documentation of ComplexityMeasures.jl.

### v3

@@ Abstract
In the nonlinear timeseries analysis literature, countless quantities have been presented as new “entropy” or “complexity” measures, often with similar roles. The ever-increasing pool of such measures makes creating a sustainable and all-encompassing software for them difficult both conceptually and pragmatically. Such a software however would be an important tool that can aid researchers make an informed decision of which measure to use and for which application, as well as accelerate novel research. Here we present ComplexityMeasures.jl, an easily extendable and highly performant open-source software that implements a vast selection of complexity measures. The software provides 1638 measures with 3,841 lines of source code, averaging only 2.3 lines of code per exported quantity (version 3.7). This is made possible by its mathematically rigorous composable design. In this paper we discuss the software design and demonstrate how it can accelerate complexity-related research in the future. We carefully compare it with alternative software and conclude that ComplexityMeasures.jl outclasses the alternatives in several objective aspects of comparison, such as computational performance, overall amount of measures, reliability, and extendability. ComplexityMeasures.jl is also a component of the DynamicalSystems.jl library for nonlinear dynamics and nonlinear timeseries analysis and follows open source development practices for creating a sustainable community of developers and contributors.
@@ 1 Introduction
A large aspect of nonlinear timeseries analysis [1, 2, 3] is concerned with extracting quantities (i.e. computing various statistics) from timeseries that quantify some property of the underlying dynamics that generated the timeseries. The purpose of these statistics can be to distinguish one type of dynamics from another [4], to classify timeseries into classes with different dynamics [5, 6], to quantify directional associations between time series [7], which in turn can be integrated into frameworks for conditional independence testing between time series  [8], and more. Most of these statistics are labeled complexity measures, because they quantify in some way the amount of complexity in the system. Although the word “complex” does not have a widely-accepted definition yet [9], it typically describes something that is both not regular nor purely stochastic. There is a multitude (100+) of software that implement some form of complexity measures. An excellent summary of some of these software is given in the supplementary material of [10].
The majority of complexity measures are based on some form of axiomatically well-founded entropy. For example, the permutation entropy [11] and the wavelet entropy [12] are based on the Shannon entropy (Eq. 1). Other complexity measures are not entropies in the formal mathematical sense, but are inspired by, or related to, entropies. Approximate entropy [13] and sample entropy [14], for example, are entropy rates, rather than entropies. Like an entropy, these entropy-like measures will typically yield higher numerical values for more “complex” data, where ”complex” has a measure-specific definition. In the rest of the text we will refer to all these entropy or complexity quantities simply as complexity measures, and only use the word “entropy” if its rigorous mathematical definition is important in the context.
@@ 1.1 Example of estimating a complexity measure
There exists a surprisingly large amount of complexity measures in the literature. For simplicity, we will start by focusing on the largest class of complexity measures: those that are functionals of probabilities mass functions (PMFs). Here, we will deal exclusively with empirical PMFs, which are PMFs estimated from data. We will use the terms “probabilities” and empirical PMFs interchangeably.
All discrete probability-based complexity measures apply the same fundamental steps, and can be unified under one estimation pipeline. To estimate probabilities from observed data, it is necessary to first define a specific and countable outcome space $\Omega$. The goal is then to assign a probability $p_{i}=p(\omega_{i})$ to each outcome $\omega_{i}\in\Omega$ based on the input data. To do so, the input data needs to be mapped (encoded, or discretized) onto the elements of $\Omega$. Sometimes, this procedure is also called symbolization. After encoding, we can form an empirical distribution over the encoded symbols. An empirical distribution in this context just means a histogram, or a normalized pseudo-histogram, depending on the structure of the outcome space. Next, the probabilities $p(\omega_{i})$ are estimated from this empirical distribution, for example using relative frequency estimation. Finally, once a probability vector $\mathbf{p}=\{p(\omega_{i})\}_{i=1}^{N}$ has been constructed, these probabilities can be given as input to some probabilities functional.
As an example, let’s say we want to compute the order-3 permutation entropy [11] for an input time series $x$. To do so, we first construct a 3-dimensional embedding of $x$. Three-element state vectors can be ordered in $3!=6$ different ways. We’ll consider each one of these possible orderings, which are also called ordinal patterns, as separate outcomes. We can then proceed by mapping each state vector in the embedding onto one of the ordinal patterns  [^1] Assuming there are no ties in the state vector., which is the same as saying that we encode the input data. To estimate probabilities, we can then simply count the relative frequency/occurrence of the different ordinal patterns and normalize these counts to sum to 1 (also called plug-in, or maximum likelihood estimation). Finally, these probabilities are given to the Shannon entropy formula [15]
@@
This estimator of the Shannon entropy is called the naive, or plug-in estimator, and returns a nonnegative number which is indicative of the "complexity" of the input data.
@@ 1.2 Combinatorial explosion of complexity measures
In the first step of the procedure outlined above, we implicitly chose an outcome space (a way to map data into outcomes). We could use any other outcome space instead. One commonly used class of outcome spaces are rectangular binnings (i.e., histograms), in which each data point is mapped onto one bin according to its value. Other examples of outcome spaces are dispersion patterns [16], binned cosine similarities [17], binned state vector distances [18], or sorting complexity [19], which all consist of an initial embedding step, after which the embedding vectors are encoded using some procedure that cleverly highlights some interesting property of the underlying data.
@@
In the third step, we could have instead considered any of the other theoretically well-founded information measures, for example Rényi entropy ($H_{R}$) [23], Tsallis entropy ($H_{T}$) [24], Kaniadakis entropy ($H_{K}$) [25], Curado entropy ($H_{C}$) [26] and the Anteneodo-Plastino stretched exponential entropy ($H_{AP}$) [27], the lesser known Shannon extropy ($J_{S}$) [28], Rényi extropy [29] or Tsallis extropy ($J_{T}$) [30], fluctuation complexity [31], or any other probabilities functional that in some way quantify complexity.
@@
Thus, excluding parameterizations, there are four degrees of freedom when computing a discrete, probabilities-based complexity measure: the discretization/encoding procedure, the probabilities estimator, the information measure, and the estimator for the information measure. Now let’s assume that the literature describes $N_{O}$ different outcome spaces, $N_{P}$ different ways of estimating probabilities, and $N_{C}$ different probabilities-based complexity measures. Assume that the number of estimators for a particular measure $M_{i}$ is $N_{M_{i}}$. Then the number of total computable PMF-based complexity measures is
@@
We quickly realize that there is a vast set possible complexity measures, varying across all four degrees of freedom, with differences ranging from minor technicalities to major conceptual ones, yet all quantifying complexity in some unique way. Adding (for example) one more entropy definition $N_{C}\to N_{C}+1$ drastically increases the total number of computable measures, because there are $N_{O}\times N_{P}$ potential ways of computing it from data for every unique estimator of this new measure. We call this the combinatorial explosion of complexity measures. In this context, the traditional software design approach of implementing one function for each measure is not scalable as it requires adding many more functions when wanting to add “only” one more complexity measure definition.
The large number of possible complexity measures could provide great opportunities for future research into complexity quantification. However, a systematic and easy-to-use software for exploring and comparing these different measures, which is also easy to extend with new ones and has the computational capacity to compute all of them quickly, is lacking.
@@
ComplexityMeasures.jl is recently developed Julia-based [40] software that fills this gap. We chose the Julia language because it is performant, high level, has an overwhelmingly open source community, and has a mature and feature-full software ecosystem that supports nonlinear timeseries analysis (§2.5). In several scientific domains authors provide similar arguments for the adoption of Julia for computational research [41, 42, 43].
ComplexityMeasures.jl resolves the explosion problem by taking a fundamentally different approach: the software allows for composable instructions for how to compute a complexity measure. For discrete, probabilities-based measures, this entails composing instructions on which definition to use, which estimator to use, with which probabilities estimator to estimate probabilities, and which outcome space to use for the discretization (Figure 1). A similar approach is taken when estimating other complexity measures, which are described later in the paper. In practice, this leads to an incredibly lean source code base that is also easily extendable. This new design approach avoids the one-function-per-estimator typical software design that leads to an unnecessarily large, hard-to-maintain code base that is also hard to bug fix efficiently or future-proof. By developing ComplexityMeasures.jl following highest standards in scientific software development [44], we not only made the software capable of handling all present and future complexity measures, but also made it highly performant.
In Sec. 2 we expose the design behind ComplexityMeasures.jl and how it integrates into a wider ecosystem enabling further innovation (such as Associations.jl  [45]); in Sec. 3 we provide some selected examples that highlight how accessible yet powerful the library is; in Sec. 4 we compare the software with existing alternatives in a comprehensive tabular format and show that across many objective aspects of comparison ComplexityMeasures.jl performs best; and lastly in Sec. 5 we first summarize our work and present a list of the novelties and unique advantages of ComplexityMeasures.jl, and conclude with an outlook on the role of open source in the wider research community, our approach on promoting community efforts, and an invite to authors developing new complexity measures.
@@ 2.1 Core functions
The design of ComplexityMeasures.jl is displayed in Fig. 1 and perfectly parallelizes the mathematically rigorous formulation of an information measure as we described it in Sec. 1.1. Software usage revolves around two functions: information and complexity. The first is called as information(discr_measure_est, prob_est, ospace, data), and estimates an information measure given the information measure estimator from a PMF, a probabilities estimator to map discrete outcomes into a PMF, and an outcome space to map input data into discrete outcomes. Internally in the software, this system is based on a functional approach that utilizes the multiple dispatch paradigm of the Julia language. When information is called many lower-level functions are called in sequence. The first function transforms the data to a symbolic representation using ospace. Multiple dispatch decides which concrete function implementation is used based on the type of ospace. Next, the symbolic representation is transformed into a PMF based on the type of prob_est. Finally, the estimated PMF is used to calculate an information measure, e.g. an entropy, based on the type of discr_measure_est. All other functions of ComplexityMeasures.jl operate in the same functional principle that utilizes multiple dispatch [^2] In practice the majority of ComplexityMeasures.jl actually utilizes single dispatch, not multiple. We refer to it as “multiple” due to it being the standard name used within the Julia programming language..
For an example of usage, to estimate the permutation entropy (as originally defined in [11]) one would call information(Shannon, RelativeAmount, OrdinalPatterns(m = 3), input). For convenience and conciseness, simpler “shortcuts” are also possible in ComplexityMeasures.jl. For example, entropy(OrdinalPatterns(m = 3), data) uses default values for the probability estimator (RelativeAmount()) and the information measure estimator (Shannon()). For a few handpicked measures even simpler syntax is available, such as the call entropy_permutation(data; m = 3), which is equivalent with the previous ones. More syntax shortcuts are discussed in the software documentation [46].
Fig 1: The two central functions of ComplexityMeasures.jl and the abstract inputs they expect, along with a couple of concrete example inputs. The full list of concrete possible inputs is displayed in Table 2. A runnable concrete code snippet that performs real-world analysis using ComplexityMeasures.jl can be seen in Appendix S1 Code snippet for Figure 3..
For complexity measures that are not explicit functionals of a PMF, we have implemented a simpler design shown in Fig. 1b. There, a function complexity takes as input a “complexity estimator” that defines a complexity measure and includes instructions on how to compute it directly from input data. Differential (or continuous) information measure estimators follow the same design in the current software version. In the future, we plan to make a dedicated probability density estimation interface for continuous measures, similar to our discrete estimation interface based on probability mass functions. Both information and complexity have a normalized option that returns the complexity measure divided by its maximum possible value. Similarly, a multiscale function computes the multiscale variant of a chosen complexity measure (a framework originally introduced in Ref. [47] as “multiscale sample entropy”). Both normalized and multiscale forms are valid for any complexity measure in the library since these forms are also based on a composable design.
@@
    Orthogonality of inputs (information estimator, probabilities, and outcomes). This means that if, e.g., a combination of a probability estimator works for a given outcome space, it is guaranteed to work for any outcome space implemented in the software. This guarantee occurs automatically due to the design, and is made possible by the Julia language multiple dispatch system [40]. It does not need to be enforced by the user or even the developer.
@@
    The software is easy to maintain. Finding a bug, for example in the process of creating a histogram of some (optionally symbolized) data, requires fixing this bug in only a single function that does the histogram counting, not in potentially hundreds of functions that utilize histograms in some way.
@@
    Simple and scalable extensions. Adding a new outcome space requires writing code for one new type (Julia’s version of “classes”) and one mandatory function extension instructing how to discretize data for this outcome space. This can be as simple as 10 lines of code. Yet, once implemented, this outcome space would allow the user, without writing any additional code, to compute any probabilities-based complexity measure using this outcome space, in combination with any probabilities estimation, any information measure, and any estimator of this measure, including even some non-information complexity measures such as missing patterns [48].
@@
    The probabilities themselves are directly accessible by the user and compose a comprehensive interface that we expand more in Sec. 2.2. Analyzing the probabilities directly may expose something interesting about the data that is “integrated away” by the complexity measure computation. For example, if one computes the associated probabilities for the 6 order-3 ordinal patterns of a timeseries coming for a logistic map, 1 of the 6 patterns will always have 0 probability due to the logistic map’s dynamics. Information like this can only be obtained by looking at the probability mass function directly, which is hidden away in the majority of alternative software. Additionally, having access to the probabilities directly allows expanding and creating new complexity measures not published before like the example we showcase in Sec. 3.2.
@@ 2.2 Outcome spaces and probabilities
The majority of complexity measures are estimated based on some PMF extracted from data. To create the extendable design mentioned above, behind the main functions information, complexity there exists a fully-fledged API (application programming interface) for extracting probabilities from data, based on the mathematically rigorous formulation of an outcome space. No other software on complexity analysis provides such an extensive interface for extracting probabilities from data.
In ComplexityMeasures.jl we define an abstract hierarchy of types called OutcomeSpace. An instance o of a concrete implementation of an OutcomeSpace describes how to discretize data into discrete outcomes. Given o, we define many functions for handling probabilities. For example, total_outcomes returns the cardinality of o, while missing_outcomes returns the number of outcomes defined by o as possible but not present in the data, and counts returns a vector of integers, counting the number of times each outcome was present in the input data. We further separate this API into outcome spaces that are counting-based, which allow mapping each element of input data into an integer, and into non-counting-based outcome spaces, which cannot do this mapping. More details on this can be found in the developer documentation of ComplexityMeasures.jl [46].